13 Question (5 points) K3L3 is a good approximation of actual An economic analysis of the production of laptop computers concSuppose that, in the short run, a firm has 27 units of capital. What can we conclude about the marginal product of labor? ChoPart 2 (1 point) See Hint Suppose that capital is still fixed at 27 units and that the prevailing wage is $45.00 while the rePart 3 (1 point) See Hint Now consider the long-run problem where capital and labor are both variable. With wages of $45.00 aPart 4 (1 point) & Feedback See Hint Based on your answers from Parts 2 and 3, what can you conclude? X Choose one or more: M
Answer #1

Flax, d.) = (dk) 3 3 24 ho Rus Increasing rehwen to scal exist. Aus is A K-27, d= (27) 3,2 - 90% MP = 20 = 3 (9) { sure Mhis